# Measuring Vision-Language STEM Skills of Neural Models

### v1

@@ Appendix D Summary of Skills
Subject: Engineer Description: iption: Which of the following could Luke’s test show? Luke had a cookie recipe that made soft, thick cookies. But he preferred crunchy cookies. Luke read that using different types of sugar affects how firm the cookies are. His recipe used both white and brown sugar, so he decided to see if the cookies would be crunchy if he didn’t use any brown sugar. Luke baked a batch of cookies using his recipe, but he left out the brown sugar and doubled the amount of white sugar. He baked the cookies for the same amount of time as in his original recipe. After the cookies finished baking and cooling, he tried one to find out how firm it was. The passage below describes how the engineering-design process was used to test a solution to a problem. Read the passage. Then answer the question below. Picture: [Uncaptioned image] Choices: [if cookies made with only white sugar were soft, if baking cookies for longer made them more crunchy, if cookies made with double the amount of brown sugar were crunchy, ] Answer index: 0
@@
Subject: Engineer Description: iption: Identify the question that Myra’s experiment can best answer. The passage below describes an experiment. Read the passage and then follow the instructions below. Myra glued lids onto 16 cardboard shoe boxes of equal size. She painted eight of the boxes black and eight of the boxes white. Myra made a small hole in the side of each box and then stuck a thermometer partially into each hole so she could measure the temperatures inside the boxes. She placed the boxes in direct sunlight in her backyard. Two hours later, she measured the temperature inside each box. Myra compared the average temperature inside the black boxes to the average temperature inside the white boxes. Picture: [Uncaptioned image] Choices: [Do the temperatures inside boxes depend on the sizes of the boxes?, Do the insides of white boxes get hotter than the insides of black boxes when the boxes are left in the sun?, ] Answer index: 1
@@
Subject: Engineer Description: iption: Which of the following could Zoe and Evelyn’s test show? Zoe and Evelyn were making batches of concrete for a construction project. To make the concrete, they mixed together dry cement powder, gravel, and water. Then, they checked if each batch was firm enough using a test called a slump test. They poured some of the fresh concrete into an upside-down metal cone. They left the concrete in the metal cone for 30 seconds. Then, they lifted the cone to see if the concrete stayed in a cone shape or if it collapsed. If the concrete in a batch collapsed, they would know the batch should not be used. The passage below describes how the engineering-design process was used to test a solution to a problem. Read the passage. Then answer the question below. Picture: [Uncaptioned image] Choices: [if the concrete from each batch took the same amount of time to dry, if a new batch of concrete was firm enough to use, ] Answer index: 1
@@
Subject: Engineer Description: iption: Identify the question that Belle’s experiment can best answer. The passage below describes an experiment. Read the passage and then follow the instructions below. Belle planted 25 tomato seeds one-half inch below the soil surface in each of six pots. Belle added an equal amount of fertilizer to three of the six pots. She placed the pots in a plant growth chamber where all the seeds experienced the same temperature, amount of light, and humidity level. After two weeks, Belle counted the number of seedlings that grew in each pot. She compared the number of seedlings in the pots with fertilizer to the number of seedlings in the pots without fertilizer. Picture: [Uncaptioned image] Choices: [Do more tomato seedlings grow when they are planted in soil with fertilizer compared to soil without fertilizer?, Does the humidity level where tomato seeds are planted affect the number of tomato seedlings that grow?, ] Answer index: 0
@@
Subject: Engineer Description: iption: In this experiment, which were part of a control group? The passage below describes an experiment. After a severe winter storm, Sandeep’s driveway was covered with ice. He read that salt makes ice melt at a lower temperature. Before covering his entire driveway with salt, he wanted to know if adding salt could actually help melt ice in the freezing outdoor temperatures. Sandeep weighed twenty ice cubes. He sprinkled salt on half of the ice cubes and left the other half unsalted. He placed all the ice cubes outside. One hour later, Sandeep quickly dried each ice cube and reweighed it to see how much it had melted. Picture: [Uncaptioned image] Choices: [the salted ice cubes, the unsalted ice cubes, ] Answer index: 1
@@
Subject: Science Description: Select the gray mineral. Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 1
Subject: Science Description: Select the one substance that is not a mineral. Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 0
@@
Subject: Science Description: iption: This organism is a spot-fin porcupinefish. Its scientific name is Diodon hystrix. Select the organism in the same genus as the spot-fin porcupinefish. Picture: [Uncaptioned image] Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 1
Subject: Science Description: Select the liquid. Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 3
@@
Subject: Science Description: iption: Fish are a group of animals with similar traits. The following traits can be used to identify fish: They have fins, not limbs. They make eggs with no shells. Observe the animals and read the descriptions. Select the one animal that has all of the fish traits listed above. Brown pelicans live along the west coast of North America. They dive underwater to catch fish in their beaks. Brown pelicans keep their eggs warm by standing on the shells with their large, webbed feet. Salmon lay eggs with no shells at the bottom of freshwater streams. Salmon use their powerful fins to swim. They can even jump up small waterfalls! Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 0
@@
Subject: Science Description: iption: Two identical blocks are heated to different temperatures. The blocks are placed so that they touch each other. Heat can flow from one block to another but cannot escape from the blocks. Later, the temperature of each block is measured again. Which pair of temperatures is possible? Picture: [Uncaptioned image] Choices: [Uncaptioned image][Uncaptioned image] Answer index: 0
@@
Subject: Science Description: iption: Two solid blocks are heated to the temperatures shown. The blocks are placed so they touch. Which diagram shows the direction heat will flow? Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 0
Subject: Science Description: Select the plant. Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 0
@@
Subject: Science Description: iption: Use the data to answer the question below. Is the following statement about our solar system true or false? Of the four smallest planets, two are made mainly of gas. Picture: [Uncaptioned image] Choices: [false, true, ] Answer index: 0
@@
Subject: Science Description: iption: Think about the magnetic force between the magnets in each pair. Which of the following statements is true? The images below show two pairs of magnets. The magnets in different pairs do not affect each other. All the magnets shown are made of the same material, but some of them are different sizes. Picture: [Uncaptioned image] Choices: [The magnitude of the magnetic force is smaller in Pair 2., The magnitude of the magnetic force is the same in both pairs., The magnitude of the magnetic force is smaller in Pair 1., ] Answer index: 0
@@
Subject: Science Description: iption: Think about the magnetic force between the magnets in each pair. Which of the following statements is true? The images below show two pairs of magnets. The magnets in different pairs do not affect each other. All the magnets shown are made of the same material. Picture: [Uncaptioned image] Choices: [The magnetic force is stronger in Pair 2., The magnetic force is stronger in Pair 1., The strength of the magnetic force is the same in both pairs., ] Answer index: 0
Subject: Science Description: Select the gas. Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 2
Subject: Science Description: Select the plant. Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 0
Subject: Science Description: iption: Which property matches this object? Select the better answer. Picture: [Uncaptioned image] Choices: [soft, smooth, ] Answer index: 1
Subject: Science Description: iption: Select the animal that does not have a backbone. Hint: Insects, spiders, and worms do not have backbones. Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 0
Subject: Science Description: iption: The diagram below is a model of two solutions. Each green ball represents one particle of solute. Which solution has a higher concentration of green particles? Picture: [Uncaptioned image] Choices: [neither; their concentrations are the same, Solution A, Solution B, ] Answer index: 2
Subject: Science Description: iption: Two solid blocks are at different temperatures. The blocks are touching. Which picture shows how heat will move? Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 1
Subject: Science Description: Select the chemical formula for this molecule. Picture: [Uncaptioned image] Choices: [H2C, HCl, HC, HCl2, ] Answer index: 1
Subject: Science Description: iption: Which statement best describes the climate of Bangor? Hint: Summers in the Northern Hemisphere occur in June, July, and August. Winters in the Northern Hemisphere occur in December, January, and February. Bangor, Maine, is a city in the United States. It has a warm summer continental climate. Picture: [Uncaptioned image] Choices: [Summers have higher temperatures and slightly more precipitation than winters., On average, On average, ] Answer index: 1
Subject: Science Description: Select the temperature shown by this thermometer. Picture: [Uncaptioned image] Choices: [13°F, 61°F, 56°F, ] Answer index: 2
@@
Subject: Math Description: iption: This table shows Jason’s January budget. What could Jason do to balance his budget? Picture: [Uncaptioned image] Choices: [increase income from shoveling snow to $40,spend$15 less at Pizza Palace, spend only \$40 at the arcade, spend \$20 more on video games, ] Answer index: 2
Subject: Math Description: In solving this triangle, which law must you use first? Picture: [Uncaptioned image] Choices: [Law of Cosines, Law of Sines, ] Answer index: 1
Subject: Math Description: Is this angle acute, right, obtuse, or straight? Picture: [Uncaptioned image] Choices: [straight, obtuse, acute, right, ] Answer index: 1
@@
Subject: Math Description: iption: Look at this cube: If the side lengths are tripled, then which of the following statements about its volume will be true? Picture: [Uncaptioned image] Choices: [The ratio of the new volume to the old volume will be 81:1., The ratio of the new volume to the old volume will be 1:8., The ratio of the new volume to the old volume will be 3:1., The ratio of the new volume to the old volume will be 27:1., ] Answer index: 3
Subject: Math Description: Which shape is a cone? Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 2
Subject: Math Description: Is the function f(x) continuous on the open interval (3,7)? Picture: [Uncaptioned image] Choices: [no, yes, ] Answer index: 1
@@
Subject: Math Description: iption: Use the diagram to help you answer the question below. Which of the following is a rational number but not an integer? Picture: [Uncaptioned image] Choices: [–123, 83, 194, 6.53, ] Answer index: 3
Subject: Math Description: Is this polygon a trapezoid? Picture: [Uncaptioned image] Choices: [no, yes, ] Answer index: 1
@@
Subject: Math Description: Look at the colored part of each shape. Which shape shows one-third? Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 0
Subject: Math Description: Which shape has 5 equal sides? Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 1
@@
Subject: Math Description: iption: Identify the cross section of this object. Assume objects are perpendicular if they appear so. Picture: [Uncaptioned image] Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 0
Subject: Math Description: Are there more circles or triangles? Picture: [Uncaptioned image] Choices: [ circles, triangles, ] Answer index: 0
@@
Subject: Math Description: iption: Look at this shape: Which image shows a reflection? Picture: [Uncaptioned image] Choices: [C, A, B, ] Answer index: 2
Subject: Math Description: Is the function f(x) continuous? Picture: [Uncaptioned image] Choices: [no, yes, ] Answer index: 0
@@
Subject: Math Description: An ice cream sundae costs 1 dollar and 41 cents. Do you have enough money to buy it? Picture: [Uncaptioned image] Choices: [yes, no, ] Answer index: 0
Subject: Math Description: Which shape has a triangle as a face? Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 1
@@
Subject: Math Description: iption: Look at this figure: What is the shape of its bases? Picture: [Uncaptioned image] Choices: [decagon, octagon, rectangle, circle, ] Answer index: 2
Subject: Math Description: The graph below shows a function. Is its inverse also a function? Picture: [Uncaptioned image] Choices: [yes, no, ] Answer index: 0
Subject: Math Description: What is the range of this exponential function? Picture: [Uncaptioned image] Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 0
Subject: Math Description: iption: Look at this graph: Is this relation a function? Picture: [Uncaptioned image] Choices: [no, yes, ] Answer index: 1
@@
Subject: Engineer Skill: evaluate-tests-of-engineering-design-solutions Description: Which of the following could Eliana’s test show?Eliana was taking part in her school’s engineering competition. To win the competition, she needed to build the popsicle-stick bridge that would hold the most weight. She could use only 200 popsicle sticks. She had two different design ideas. She had to pick one of the designs to use in the competition. To test which design was strongest, Eliana built two prototypes, each with 200 popsicle sticks. She then added 1 kg weights to each prototype until one of them broke. The passage below describes how the engineering-design process was used to test a solution to a problem. Read the passage. Then answer the question below. Picture: [Uncaptioned image] Choices: [how much weight a bridge built with 300 popsicle sticks could hold, which design could hold more weight, ] Answer index: 1
@@
Subject: Engineer Skill: identify-control-and-experimental-groups Description: In this experiment, which were part of a control group?The passage below describes an experiment. Madelyn has a bubble machine and wants to know how to make the bubbles last longer. She read that bubbles burst when the liquid that makes up the bubbles evaporates. Madelyn knew that when liquids are warmer, they evaporate faster. So, she wondered if she could make her bubbles last longer by cooling the bubble solution. Madelyn cooled six bottles of bubble solution to 30°F below room temperature. She left another six bottles of bubble solution at room temperature. Then, she measured how long bubbles made from the solution in each bottle lasted. Picture: [Uncaptioned image] Choices: [the bottles that were cooled down, the bottles that were at room temperature, ] Answer index: 1
@@
Subject: Engineer Skill: identify-independent-and-dependent-variables Description: Hint: An independent variable is a variable whose effect you are investigating. A dependent variable is a variable that you measure.Which of the following was a dependent variable in this experiment? The passage below describes an experiment. Read the passage and think about the variables that are described. Giardia is a microscopic parasite that lives in water and can infect humans. Dr. Roth designed a drinking straw that contained a filter to remove Giardia from water. Dr. Roth wanted to know if a longer filtering straw would remove more Giardia. Dr. Roth made six filtering straws: three that were five inches long and three that were ten inches long. She prepared six one-liter batches of water, each containing 10,000 Giardia. Then, Dr. Roth passed one batch of water through each straw. After each batch passed through the straw, she used a microscope to count the number of Giardia that remained in a small sample of the water. Picture: [Uncaptioned image] Choices: [the number of Giardia that remained in the water, the length of the filtering straw, ] Answer index: 0
Subject: Engineer Skill: identify-laboratory-tools Description: Select the round-bottom flask. Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 1
@@
Subject: Engineer Skill: identify-the-experimental-question Description: Identify the question that Jeffrey’s experiment can best answer.The passage below describes an experiment. Read the passage and then follow the instructions below. Jeffrey mixed bacteria into a nutrient-rich liquid where the bacteria could grow. He poured four ounces of the mixture into each of ten glass flasks. In five of the ten flasks, he also added one teaspoon of cinnamon. He allowed the bacteria in the flasks to grow overnight in a 37°C room. Then, Jeffrey used a microscope to count the number of bacteria in a small sample from each flask. He compared the amount of bacteria in the liquid with cinnamon to the amount of bacteria in the liquid without cinnamon. Picture: [Uncaptioned image] Choices: [Do more bacteria grow in liquid with cinnamon than in liquid without cinnamon?, Does temperature affect how much bacteria can grow in liquid?, ] Answer index: 0
Subject: Engineer Skill: laboratory-safety-equipment Description: Select the apron. Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 3

### v2

@@ Appendix D Summary of Skills
Description: iption: Which of the following could Luke’s test show? Luke had a cookie recipe that made soft, thick cookies. But he preferred crunchy cookies. Luke read that using different types of sugar affects how firm the cookies are. His recipe used both white and brown sugar, so he decided to see if the cookies would be crunchy if he didn’t use any brown sugar. Luke baked a batch of cookies using his recipe, but he left out the brown sugar and doubled the amount of white sugar. He baked the cookies for the same amount of time as in his original recipe. After the cookies finished baking and cooling, he tried one to find out how firm it was. The passage below describes how the engineering-design process was used to test a solution to a problem. Read the passage. Then answer the question below.
Picture: [Uncaptioned image]
Choices: [if cookies made with only white sugar were soft, if baking cookies for longer made them more crunchy, if cookies made with double the amount of brown sugar were crunchy, ]
Answer index: 0
Subject: Engineer
@@
Description: iption: Identify the question that Myra’s experiment can best answer. The passage below describes an experiment. Read the passage and then follow the instructions below. Myra glued lids onto 16 cardboard shoe boxes of equal size. She painted eight of the boxes black and eight of the boxes white. Myra made a small hole in the side of each box and then stuck a thermometer partially into each hole so she could measure the temperatures inside the boxes. She placed the boxes in direct sunlight in her backyard. Two hours later, she measured the temperature inside each box. Myra compared the average temperature inside the black boxes to the average temperature inside the white boxes.
Picture: [Uncaptioned image]
Choices: [Do the temperatures inside boxes depend on the sizes of the boxes?, Do the insides of white boxes get hotter than the insides of black boxes when the boxes are left in the sun?, ]
Answer index: 1
Subject: Engineer
@@
Description: iption: Which of the following could Zoe and Evelyn’s test show? Zoe and Evelyn were making batches of concrete for a construction project. To make the concrete, they mixed together dry cement powder, gravel, and water. Then, they checked if each batch was firm enough using a test called a slump test. They poured some of the fresh concrete into an upside-down metal cone. They left the concrete in the metal cone for 30 seconds. Then, they lifted the cone to see if the concrete stayed in a cone shape or if it collapsed. If the concrete in a batch collapsed, they would know the batch should not be used. The passage below describes how the engineering-design process was used to test a solution to a problem. Read the passage. Then answer the question below.
Picture: [Uncaptioned image]
Choices: [if the concrete from each batch took the same amount of time to dry, if a new batch of concrete was firm enough to use, ]
Answer index: 1
Subject: Engineer
@@
Description: iption: Identify the question that Belle’s experiment can best answer. The passage below describes an experiment. Read the passage and then follow the instructions below. Belle planted 25 tomato seeds one-half inch below the soil surface in each of six pots. Belle added an equal amount of fertilizer to three of the six pots. She placed the pots in a plant growth chamber where all the seeds experienced the same temperature, amount of light, and humidity level. After two weeks, Belle counted the number of seedlings that grew in each pot. She compared the number of seedlings in the pots with fertilizer to the number of seedlings in the pots without fertilizer.
Picture: [Uncaptioned image]
Choices: [Do more tomato seedlings grow when they are planted in soil with fertilizer compared to soil without fertilizer?, Does the humidity level where tomato seeds are planted affect the number of tomato seedlings that grow?, ]
Answer index: 0
Subject: Engineer
@@
Description: iption: In this experiment, which were part of a control group? The passage below describes an experiment. After a severe winter storm, Sandeep’s driveway was covered with ice. He read that salt makes ice melt at a lower temperature. Before covering his entire driveway with salt, he wanted to know if adding salt could actually help melt ice in the freezing outdoor temperatures. Sandeep weighed twenty ice cubes. He sprinkled salt on half of the ice cubes and left the other half unsalted. He placed all the ice cubes outside. One hour later, Sandeep quickly dried each ice cube and reweighed it to see how much it had melted.
Picture: [Uncaptioned image]
Choices: [the salted ice cubes, the unsalted ice cubes, ]
Answer index: 1
@@
Subject: Science
Description: Select the gray mineral.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 1
Subject: Science
Description: Select the one substance that is not a mineral.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Science
@@
Description: iption: This organism is a spot-fin porcupinefish. Its scientific name is Diodon hystrix. Select the organism in the same genus as the spot-fin porcupinefish.
Picture: [Uncaptioned image]
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 1
Subject: Science
Description: Select the liquid.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 3
Subject: Science
@@
Description: iption: Fish are a group of animals with similar traits. The following traits can be used to identify fish: They have fins, not limbs. They make eggs with no shells. Observe the animals and read the descriptions. Select the one animal that has all of the fish traits listed above. Brown pelicans live along the west coast of North America. They dive underwater to catch fish in their beaks. Brown pelicans keep their eggs warm by standing on the shells with their large, webbed feet. Salmon lay eggs with no shells at the bottom of freshwater streams. Salmon use their powerful fins to swim. They can even jump up small waterfalls!
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Science
@@
Description: iption: Two identical blocks are heated to different temperatures. The blocks are placed so that they touch each other. Heat can flow from one block to another but cannot escape from the blocks. Later, the temperature of each block is measured again. Which pair of temperatures is possible?
Picture: [Uncaptioned image]
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Science
@@
Description: iption: Two solid blocks are heated to the temperatures shown. The blocks are placed so they touch. Which diagram shows the direction heat will flow?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Science
Description: Select the plant.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Science
@@
Description: iption: Use the data to answer the question below. Is the following statement about our solar system true or false? Of the four smallest planets, two are made mainly of gas.
Picture: [Uncaptioned image]
Choices: [false, true, ]
Answer index: 0
Subject: Science
@@
Description: iption: Think about the magnetic force between the magnets in each pair. Which of the following statements is true? The images below show two pairs of magnets. The magnets in different pairs do not affect each other. All the magnets shown are made of the same material, but some of them are different sizes.
Picture: [Uncaptioned image]
Choices: [The magnitude of the magnetic force is smaller in Pair 2., The magnitude of the magnetic force is the same in both pairs., The magnitude of the magnetic force is smaller in Pair 1., ]
Answer index: 0
@@
Subject: Science
Description: iption: Think about the magnetic force between the magnets in each pair. Which of the following statements is true? The images below show two pairs of magnets. The magnets in different pairs do not affect each other. All the magnets shown are made of the same material.
Picture: [Uncaptioned image]
Choices: [The magnetic force is stronger in Pair 2., The magnetic force is stronger in Pair 1., The strength of the magnetic force is the same in both pairs., ]
Answer index: 0
Subject: Science
Description: Select the gas.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 2
Subject: Science
Description: Select the plant.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Science
Description: iption: Which property matches this object? Select the better answer.
Picture: [Uncaptioned image]
Choices: [soft, smooth, ]
Answer index: 1
Subject: Science
Description: iption: Select the animal that does not have a backbone. Hint: Insects, spiders, and worms do not have backbones.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Science
Description: iption: The diagram below is a model of two solutions. Each green ball represents one particle of solute. Which solution has a higher concentration of green particles?
Picture: [Uncaptioned image]
Choices: [neither; their concentrations are the same, Solution A, Solution B, ]
Answer index: 2
Subject: Science
Description: iption: Two solid blocks are at different temperatures. The blocks are touching. Which picture shows how heat will move?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 1
Subject: Science
Description: Select the chemical formula for this molecule.
Picture: [Uncaptioned image]
Choices: [H2C, HCl, HC, HCl2, ]
Answer index: 1
Subject: Science
Description: iption: Which statement best describes the climate of Bangor? Hint: Summers in the Northern Hemisphere occur in June, July, and August. Winters in the Northern Hemisphere occur in December, January, and February. Bangor, Maine, is a city in the United States. It has a warm summer continental climate.
Picture: [Uncaptioned image]
Choices: [Summers have higher temperatures and slightly more precipitation than winters., On average, On average, ]
Answer index: 1
Subject: Science
Description: Select the temperature shown by this thermometer.
Picture: [Uncaptioned image]
Choices: [13°F, 61°F, 56°F, ]
Answer index: 2
@@
Subject: Math
Description: iption: This table shows Jason’s January budget. What could Jason do to balance his budget?
Picture: [Uncaptioned image]
@@
Choices: [increase income from shoveling snow to $40,spend$15 less at Pizza Palace, spend only \$40 at the arcade, spend \$20 more on video games, ]
Answer index: 2
Subject: Math
Description: In solving this triangle, which law must you use first?
Picture: [Uncaptioned image]
Choices: [Law of Cosines, Law of Sines, ]
Answer index: 1
Subject: Math
Description: Is this angle acute, right, obtuse, or straight?
Picture: [Uncaptioned image]
Choices: [straight, obtuse, acute, right, ]
Answer index: 1
Subject: Math
@@
Description: iption: Look at this cube: If the side lengths are tripled, then which of the following statements about its volume will be true?
Picture: [Uncaptioned image]
Choices: [The ratio of the new volume to the old volume will be 81:1., The ratio of the new volume to the old volume will be 1:8., The ratio of the new volume to the old volume will be 3:1., The ratio of the new volume to the old volume will be 27:1., ]
Answer index: 3
Subject: Math
Description: Which shape is a cone?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 2
Subject: Math
Description: Is the function f(x) continuous on the open interval (3,7)?
Picture: [Uncaptioned image]
Choices: [no, yes, ]
Answer index: 1
Subject: Math
@@
Description: iption: Use the diagram to help you answer the question below. Which of the following is a rational number but not an integer?
Picture: [Uncaptioned image]
Choices: [–123, 83, 194, 6.53, ]
Answer index: 3
Subject: Math
Description: Is this polygon a trapezoid?
Picture: [Uncaptioned image]
Choices: [no, yes, ]
Answer index: 1
Subject: Math
@@
Description: Look at the colored part of each shape. Which shape shows one-third?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Math
Description: Which shape has 5 equal sides?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 1
Subject: Math
@@
Description: iption: Identify the cross section of this object. Assume objects are perpendicular if they appear so.
Picture: [Uncaptioned image]
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Math
Description: Are there more circles or triangles?
Picture: [Uncaptioned image]
Choices: [ circles, triangles, ]
Answer index: 0
Subject: Math
@@
Description: iption: Look at this shape: Which image shows a reflection?
Picture: [Uncaptioned image]
Choices: [C, A, B, ]
Answer index: 2
Subject: Math
Description: Is the function f(x) continuous?
Picture: [Uncaptioned image]
Choices: [no, yes, ]
Answer index: 0
Subject: Math
@@
Description: An ice cream sundae costs 1 dollar and 41 cents. Do you have enough money to buy it?
Picture: [Uncaptioned image]
Choices: [yes, no, ]
Answer index: 0
Subject: Math
Description: Which shape has a triangle as a face?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 1
Subject: Math
@@
Description: iption: Look at this figure: What is the shape of its bases?
Picture: [Uncaptioned image]
Choices: [decagon, octagon, rectangle, circle, ]
Answer index: 2
Subject: Math
Description: The graph below shows a function. Is its inverse also a function?
Picture: [Uncaptioned image]
Choices: [yes, no, ]
Answer index: 0
Subject: Math
Description: What is the range of this exponential function?
Picture: [Uncaptioned image]
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Math
Description: iption: Look at this graph: Is this relation a function?
Picture: [Uncaptioned image]
Choices: [no, yes, ]
Answer index: 1
@@
Subject: Engineer
Skill: evaluate-tests-of-engineering-design-solutions
@@
Description: Which of the following could Eliana’s test show?Eliana was taking part in her school’s engineering competition. To win the competition, she needed to build the popsicle-stick bridge that would hold the most weight. She could use only 200 popsicle sticks. She had two different design ideas. She had to pick one of the designs to use in the competition. To test which design was strongest, Eliana built two prototypes, each with 200 popsicle sticks. She then added 1 kg weights to each prototype until one of them broke. The passage below describes how the engineering-design process was used to test a solution to a problem. Read the passage. Then answer the question below.
Picture: [Uncaptioned image]
Choices: [how much weight a bridge built with 300 popsicle sticks could hold, which design could hold more weight, ]
Answer index: 1
Subject: Engineer
Skill: identify-control-and-experimental-groups
@@
Description: In this experiment, which were part of a control group?The passage below describes an experiment. Madelyn has a bubble machine and wants to know how to make the bubbles last longer. She read that bubbles burst when the liquid that makes up the bubbles evaporates. Madelyn knew that when liquids are warmer, they evaporate faster. So, she wondered if she could make her bubbles last longer by cooling the bubble solution. Madelyn cooled six bottles of bubble solution to 30°F below room temperature. She left another six bottles of bubble solution at room temperature. Then, she measured how long bubbles made from the solution in each bottle lasted.
Picture: [Uncaptioned image]
Choices: [the bottles that were cooled down, the bottles that were at room temperature, ]
Answer index: 1
Subject: Engineer
Skill: identify-independent-and-dependent-variables
@@
Description: Hint: An independent variable is a variable whose effect you are investigating. A dependent variable is a variable that you measure.Which of the following was a dependent variable in this experiment? The passage below describes an experiment. Read the passage and think about the variables that are described. Giardia is a microscopic parasite that lives in water and can infect humans. Dr. Roth designed a drinking straw that contained a filter to remove Giardia from water. Dr. Roth wanted to know if a longer filtering straw would remove more Giardia. Dr. Roth made six filtering straws: three that were five inches long and three that were ten inches long. She prepared six one-liter batches of water, each containing 10,000 Giardia. Then, Dr. Roth passed one batch of water through each straw. After each batch passed through the straw, she used a microscope to count the number of Giardia that remained in a small sample of the water.
Picture: [Uncaptioned image]
Choices: [the number of Giardia that remained in the water, the length of the filtering straw, ]
Answer index: 0
Subject: Engineer
Skill: identify-laboratory-tools
Description: Select the round-bottom flask.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 1
Subject: Engineer
Skill: identify-the-experimental-question
@@
Description: Identify the question that Jeffrey’s experiment can best answer.The passage below describes an experiment. Read the passage and then follow the instructions below. Jeffrey mixed bacteria into a nutrient-rich liquid where the bacteria could grow. He poured four ounces of the mixture into each of ten glass flasks. In five of the ten flasks, he also added one teaspoon of cinnamon. He allowed the bacteria in the flasks to grow overnight in a 37°C room. Then, Jeffrey used a microscope to count the number of bacteria in a small sample from each flask. He compared the amount of bacteria in the liquid with cinnamon to the amount of bacteria in the liquid without cinnamon.
Picture: [Uncaptioned image]
Choices: [Do more bacteria grow in liquid with cinnamon than in liquid without cinnamon?, Does temperature affect how much bacteria can grow in liquid?, ]
Answer index: 0
Subject: Engineer
Skill: laboratory-safety-equipment
Description: Select the apron.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 3
Subject: Math
Skill: absolute-value-and-integers-word-problems